OpenStudy (shamil98):

Triple Integral at the bottom of the mess of replies. How would I go about solving it?

OpenStudy (shadowlegendx):

o_O

OpenStudy (shamil98):

\[f(x) = \frac{ 1 }{ 2 }x^4\] \[f'(x) = \frac{ 1 }{ 2 }4x^3\] \[f'(x) = 2x^3\] slope of line tangent to 1/2x^4 at (2,8). so now i must turn my sleepy brain on. ..

OpenStudy (shadowlegendx):

Function of x?

OpenStudy (shamil98):

no i did that wrong oops

OpenStudy (isaiah.feynman):

@shamil98 We solved a question like this for that unresponsive guy.

OpenStudy (shamil98):

oh no i didnt, i should sleep.

OpenStudy (shamil98):

yes yes, i know i've done this before just reviewing

Parth (parthkohli):

now plug in 2 instead of x in that f'(x)

OpenStudy (shamil98):

yep.

OpenStudy (isaiah.feynman):

@ParthKohli I like that dog you used.

Parth (parthkohli):

@Isaiah.Feynman in the end of ozymandias? that was vince's idea.

OpenStudy (isaiah.feynman):

Vince's idea was good. @ParthKohli I ORDER YOU TO USE IT NOW!

Parth (parthkohli):

I was referencing Breaking Bad. Anyway...

OpenStudy (shamil98):

oh got it . 16

OpenStudy (shamil98):

maybe 1 am is a bad time for this stuff.

OpenStudy (isaiah.feynman):

@shamil98 LOL

OpenStudy (isaiah.feynman):

I'm studying statistics its 2 pm

OpenStudy (isaiah.feynman):

Batman just keeps observing!

OpenStudy (anonymous):

I'm on brilliant.org, just solved a level 5 problemmmmm :P

Parth (parthkohli):

Are you in Pak, @Isaiah.Feynman

OpenStudy (anonymous):

get on my level sham

OpenStudy (anonymous):

noob <3

OpenStudy (isaiah.feynman):

Nope. I'm in the northern west.

Parth (parthkohli):

I did too!

OpenStudy (shamil98):

are between curves y = x^2 and y = x^3 from 1 <= x <= 3 \[\int\limits_{1}^{3} x^2 + \int\limits_{1}^{3} x^3\] x^3/3 )from 1 to 3.. x^4/4 from 1 to 3 3^3/3 - 1^3/3 3^4/4 - 1^4/4 27/3 - 1 /3 81/4 - 1/4 26/3 + 80/4 = 86/3 i did something wrong what did i do wrong.

OpenStudy (shamil98):

batman i'm on lvl 3 atm

OpenStudy (anonymous):

I was doing algebra, I just finished it, pretty easy tbh.

OpenStudy (shamil98):

oh wait i got it

OpenStudy (isaiah.feynman):

Are you solving for area?

Parth (parthkohli):

why adding the areas?

OpenStudy (shamil98):

because parth i am dumb \[\int\limits_{1}^{3} x^2 - x^3 ~dx\]

OpenStudy (anonymous):

Lmao sham

OpenStudy (shamil98):

x^3/3 - x^4/4 (3^3/3 - 3^4/4) - ( 1^3/3 - 1^4/4) 9 - 81/4 - 1/3 + 1/4 9 - 1/3 + 20 29 - 1/3 86/3

OpenStudy (shamil98):

wait i got same answer what

OpenStudy (shamil98):

oh arithmetic error

OpenStudy (shamil98):

-34/3

OpenStudy (shamil98):

Darn, it says that is wrong too.

Parth (parthkohli):

how can area b/w two curves be negative?

OpenStudy (anonymous):

You can't have a negative area sham

OpenStudy (shamil98):

nvm i made another error it was x^3 - x^2 I was looking at the graphs from x = 0 to x= 1 .... its 11.3333

OpenStudy (shamil98):

Do my failures amuse you all?

OpenStudy (anonymous):

I like that you're persistent! :)

OpenStudy (shamil98):

How do i do this?.. \[\large \int\limits_{}^{} \int\limits_{}^{} \int\limits_{B}^{} xyz^2 dV\] where B is the cuboid region bounded by the regions 0 ≤ x ≤ 1, - 1 ≤ y ≤ 2 and 0 ≤ z ≤ 3?

OpenStudy (anonymous):

Lol that's on the website right?

OpenStudy (shamil98):

yah

OpenStudy (shamil98):

level four

OpenStudy (anonymous):

I looked at it, and clicked X.

OpenStudy (anonymous):

they require work

OpenStudy (shamil98):

how do you do it though, the dV is a bit offputting to me >.>

OpenStudy (shamil98):

if it was something like dzdydx i mean that would be pretty simple

OpenStudy (shamil98):

you guys have any ideas? i'm going to head to sleep now, got school in a few hours.. but i would appreciate any thoughts on the triple integral, thanks.

OpenStudy (anonymous):

=27/4

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

Was that the full problem?

OpenStudy (anonymous):